OpenStudy (anonymous):

Solve the equations using properties of logs: a.) 3log[11](x) = 2log[11]8 b.) 2log[5](x) = 3log[5]9

OpenStudy (anonymous):

\[3\log_{11}x =2\log_{11}8, x^{3}=8^{2} x=4. 2\log_{5}x=3\log_{5}9 x^{2}=9^{3} x=27\] could you write those question again by equation program in open study? I'm not sure you meant this

OpenStudy (anonymous):

there's comma between \[3\log_{5}9 and x^{2}\]

OpenStudy (anonymous):

\[3\log_{11}(x)=2\log_{11}8 \]

OpenStudy (anonymous):

\[B.) 2\log_{5}(x)=3\log_{5}9 \]

OpenStudy (anonymous):

a)\[\log_{11} (x^3)]=[\log_{11}(8^2) \] so x^3=8^2 x=8^(2/3) x=4 b) similarly x^2=9^3 x= 27

OpenStudy (anonymous):

log11(x3)]=[log11(82) or x^3=8^2 x=8^(2/3) or x=4 another is x=27